you inherit $13000.00 and decide to invest the money in two different investments; one paying 10% and the other 14%. a year later they earn $1580. how much do you originally invewst in each account.
Answer by checkley77(12569)

.14x+.10(13,000-x)=1,580
.14x+1,300-.10x=1,580
.04x=1,580-1,300
.04x=280
x=280/.04
x=7,000 amount invested @ 14%.
13,000-7,000=6,000 amount invested @ 10%.
Proof:
.14*7,000+.10*6,000=1,580
980+600=1,580
1,580=1,580